# Centaur: a foundation model of human cognition

### v2

@@ Discussion
The present paper introduced Centaur, the first real candidate for a unified model of human cognition. We obtained Centaur by finetuning a state-of-the-art language model on Psych-101 – a novel, large-scale data set of human behavior. This approach allowed us to leverage the vast knowledge embedded in large language models, while, at the same time, aligning them with human behavior [15]. Centaur successfully captures human behavior and passes a wide range of out-of-distribution checks. It generalizes not only to unseen participants but also to new cover stories, structural variations, and entirely novel domains. In addition to analyzing the model on a behavioral level, we also conducted a series of analyses on its internal representations, in which we found an increased alignment with human neural activity.
@@
With his call for unified theories of cognition [2], Newell also outlined a set of criteria that a unified computational model should fulfill [36]. We believe that Centaur is the first model to satisfy the majority of these criteria (see Table 1). Most importantly, it (1) behaves as an almost arbitrary function of the environment, (2) operates in real time, and (3) relies on vast amounts of knowledge about the world. We provide an extended discussion on Newell’s criteria for a theory of cognition in the Supplementary Information.
@@
Table 1: Newell test for a theory of cognition. We provide an extended discussion on these criteria in the Supplementary Information.
A computational model like Centaur that can simulate and predict human behavior in any domain offers many direct applications. It may, for instance, be used for in-silico prototyping of experimental studies [37]. In this context, one could use the model to figure out which designs lead to the largest effect sizes, how to design a study to reduce the number of required participants, or to estimate the power of an effect. Thinking one step further, Centaur finds applications in the context of automated cognitive science [38, 39]. For example, it can be integrated into frameworks that utilize predictive models to guide the development of psychological theories, such as scientific regret minimization [40]. Traditionally, this requires collecting huge data sets that are used to build predictive models for the experiment at hand. Centaur removes this constraint and thereby broadens the scope of these frameworks.
The ultimate goal of the research program outlined in this paper is to leverage models like Centaur to gain deeper insights into human cognition. While the present paper takes initial steps in that direction, it also opens up exciting new avenues for future exploration. First, one could further probe Centaur’s internal representations to understand how it represents knowledge and processes information. The resulting insights could, in turn, be used to generate new hypotheses about knowledge representation and information processing in humans that could be validated in future experimental studies. We believe that tools like sparse autoencoders [41] and attention map visualization [42] provide promising avenues towards accomplishing this goal and hope to explore them in future studies.
In addition, it might also be possible to train models with different architectures from scratch using the data set that we have created in the process of this paper. Doing so would enable us to investigate the neural architecture of human cognition at a scale that could not have been done before. We might, for example, ask questions such as whether human information processing is better described by attention-based architectures [43] or architectures with a vector-based memory [44], or how much we can improve by incorporating theories from the neuroscience literature [45, 46].
Even though Psych-101 is already the broadest and largest data set of human behavior out there, we view its development as an ongoing process and plan to further develop it. The focus in its current state is largely on learning and decision-making, but eventually, we intend to include additional domains such as psycholinguistics, social psychology, and economic games. Experiments with information about individual differences are another source of neglected data in the current iteration of Psych-101. Ideally, we want to include all types of relevant information about subjects (e.g., age, personality traits, or socioeconomic status) in the prompt, such that a model trained on this data can capture individual differences. Experiments from developmental psychology or computational psychiatry provide an ideal source for this purpose. Finally, while we have already included some cross-cultural and meta-studies [47, 48, 49, 50], the current iteration has still a strong bias toward a Western, educated, industrialized, rich, and democratic (WEIRD) population [51]. To address all of these shortcomings, we have created an open-source repository and invite everyone to contribute to the next iteration of Psych-101 in an open research collaboration (https://github.com/marcelbinz/Psych-201). The goal of this effort is to provide psychological data in a standardized format that facilitates benchmarking, thereby complementing existing efforts from the neuroscience community [52, 53, 54, 55].
Conclusion
[…]
Methods
Data collection
Finetuning procedure
We used Llama 3.1 70B as the base model for our finetuning procedure. We relied on a parameter-efficient finetuning technique known as QLoRA [17]. QLoRA adds so-called low-rank adapters to each layer of a 4-bit quantized base model. The base model was kept fixed during finetuning and only the parameters of the low-rank adapters were adjusted. For further details on this technique, we refer the reader to [17]. We finetuned the model for one epoch on the entire data set using a standard cross-entropy loss (we experimented with prolonged training but found that this led to overfitting). We only backpropagated the loss at human responses and masked out the loss for all other tokens. The effective batch size was set to 32, the learning rate to 0.00005, and the weight decay to 0.01. We used an 8-bit AdamW optimizer [57] with a linearly increasing warm up over the first 100 gradient steps. We added low-rank adapters to all non-embedding layers and set both the rank as well as the scaling parameter to 8. The finetuning procedure was implemented using the unsloth library (https://unsloth.ai/).
Evaluation metric
We used a pseudo-R2 measure to evaluate all our models [18]. This measure takes the average log-likelihood of human responses for a given model and normalizes it using the average log-likelihood of a model that guesses responses uniformly:
$\text{R}^{2}=1-\dfrac{\log p_{\text{model}}(\mathcal{D})}{\log p_{\text{guess}}(\mathcal{D})}$ (1)
A value of zero corresponds to prediction at chance level while a value of one corresponds to perfect predictability. Therefore, this measure ensures a common scale across all experiments. For experiments with single-token responses, $\log p_{\text{model}}(\mathcal{D})$ simply amounted to the average log-likelihood of tokens. For experiments with multi-token responses, we summed log-likelihoods within a response and averaged across responses.
Domain-specific cognitive models
For our main analysis, we were interested in predicting the behavior of held-out participants. Therefore, we fitted a joint set of parameters for all participants in the training data and evaluated how well a model with these parameters predicts responses of held-out participants. Mirroring the evaluation metric for the language-based models, we evaluated goodness-of-fit using a predictive pseudo-R2 measure.
Neural alignment
The neural alignment analysis on the two-step task was conducted using data collected in a previous study [33]. We used a regularized linear regression model to predict fMRI data from internal representations of Centaur and Llama (a separate model was used for each participant and region). We fitted each of these models on data from two scanning blocks and evaluated them on data from the third. The regularization strength was selected using a nested cross-validation procedure. For each run, we split the beta maps into cortical and subcortical regions of interest (ROI) using the Harvard-Oxford atlas. We averaged the betas within each ROI, reducing the number of betas from the number of voxels to the number of ROIs. All cortical and subcortical ROIs from the atlas were evaluated. Reported R2 values correspond to the average across all ROIs.
Data availability
Code availability
[…]
Acknowledgements
[…]
Author contributions
Project lead: Marcel Binz
Data curation: Elif Akata, Franziska Brändle, Marcel Binz, Fred Callaway, Julian Coda-Forno, Can Demircan, Maria Eckstein, Noemi Elteto, Susanne Haridi, Akshay Jagadish, Li Ji-An, Alexander Kipnis, Sreejan Kumar, Tobias Ludwig, Surabhi Nath, Joshua Peterson, Evan Russek, Tankred Saanum, Natalia Scharfenberg, Johannes Schubert, Luca Schulze Buschoff, Nishad Singhi, Xin Sui, Mirko Thalmann, Vuong Truong, Kristin Witte, Shuchen Wu, Dirk Wulff, Huadong Xiong
Model training: Marcel Binz, Vishaal Udandarao
Model evaluation: Marcel Binz, Julian Coda-Forno, Alexander Kipnis, Konstantinos Voudouris
First draft: Marcel Binz, Eric Schulz
Review and editing: All authors
No researchers at Google DeepMind used Llama for this research.
Supplementary Information
Full goodness-of-fit results
Experiment Centaur Llama Cognitive model
Shepard categorization 0.2315 0.1695 0.0765
Drifting four-armed bandit 0.4896 0.3622 0.3225
N-back 0.4282 0.2484 0.1651
Digit span 0.8431 0.8120 0.7388
Go/no-go 1.0000 0.9910 0.8908
Recent probes 0.6274 0.5026 0.4420
Horizon task 0.4464 0.2730 0.1172
Gardening task 0.4542 0.2729 0.1509
Columbia card task 0.7063 0.6460 0.6219
Balloon analog risk task 0.9048 0.8808 0.8670
Two-armed bandit 0.5725 0.4475 0.4108
Conditional associative learning 0.5326 0.4427 0.3115
THINGS odd-one-out 0.2853 -0.0611 0.2406
Multi-attribute decision-making 0.9107 0.7833 0.7227
Two-step task 0.2715 0.1275 nan
Probabilistic instrumental learning 0.2878 0.2236 0.2802
Medin categorization 0.2835 0.1672 0.2303
Zoopermarket 0.3002 0.1306 0.1256
choices13k 0.3770 0.2075 0.0530
CPC18 0.5109 0.4058 0.0466
Intertemporal choice 0.3677 -0.0656 0.0492
Structured bandit 0.6918 0.6098 0.5136
Weather prediction task 0.2045 0.1705 0.0959
Iowa gambling task 0.3499 0.2781 0.1497
Virtual subway network 0.6864 0.5663 nan
Multi-task reinforcement learning 0.4837 0.3989 0.0625
Serial reaction time task 0.9521 0.9470 0.9453
Decisions from description 0.1478 -0.3177 0.1183
Decisions from experience 0.3236 0.0626 0.2134
Changing bandit 0.5635 0.4483 0.3685
Multiple-cue judgment 0.4888 0.4167 0.1245
Recall and recognition 0.7046 0.6162 nan
Experiential-symbolic task 0.3459 -0.0070 nan
Grammar judgement 0.2877 0.0101 0.2989
Risky choice 0.3828 0.0663 nan
Tile-revealing task 0.5188 0.2966 nan
Episodic long-term memory 0.7578 0.6837 nan
Aversive learning 0.1156 -0.1087 nan
Spatially correlated multi-armed bandit 0.4613 0.2804 0.1789
Probabilistic reasoning 0.3378 0.2632 nan
Table 2: Full pseudo-R2 results on held-out participants.
Experiment Centaur Llama Cognitive model
Shepard categorization 0.5327 0.5756 0.6401
Drifting four-armed bandit 0.7075 0.8842 0.9393
N-back 0.3963 0.5210 0.5787
Digit span 0.5622 0.6735 0.9359
Go/no-go 0.0000 0.0062 0.0757
Recent probes 0.2583 0.3448 0.3868
Horizon task 0.3837 0.5039 0.6119
Gardening task 0.3783 0.5040 0.5885
Columbia card task 0.2036 0.2454 0.2621
Balloon analog risk task 0.0660 0.0826 0.0922
Two-armed bandit 0.2963 0.3829 0.4084
Conditional associative learning 0.5135 0.6123 0.7564
THINGS odd-one-out 0.7852 1.1657 0.8343
Multi-attribute decision-making 0.0619 0.1502 0.1922
Two-step task 0.5050 0.6048 nan
Probabilistic instrumental learning 0.4937 0.5382 0.4989
Medin categorization 0.4967 0.5772 0.5335
Zoopermarket 0.4850 0.6026 0.6061
choices13k 0.4318 0.5493 0.6564
CPC18 0.3390 0.4118 0.6609
Intertemporal choice 0.4383 0.7386 0.6590
Structured bandit 0.6410 0.8114 1.0114
Weather prediction task 0.5514 0.5749 0.6267
Iowa gambling task 0.9012 1.0007 1.1787
Virtual subway network 1.1237 1.5540 nan
Multi-task reinforcement learning 0.5672 0.6604 1.0299
Serial reaction time task 0.1718 0.1900 0.1962
Decisions from description 0.5907 0.9133 0.6111
Decisions from experience 0.4688 0.6497 0.5452
Changing bandit 0.3025 0.3824 0.4378
Multiple-cue judgment 1.1232 1.2815 1.9237
Recall and recognition 1.0587 1.3754 nan
Experiential-symbolic task 0.4534 0.6980 nan
Grammar judgement 1.4351 1.9945 1.4127
Risky choice 0.4278 0.6472 nan
Tile-revealing task 1.8728 2.7374 nan
Episodic long-term memory 0.8680 1.1335 nan
Aversive learning 4.0729 5.1057 nan
Spatially correlated multi-armed bandit 1.8321 2.4475 2.7928
Probabilistic reasoning 2.3730 2.6405 nan
Table 3: Full negative log-likelihoods results on held-out participants.
Regression results
[…]
Figure 5: Regression analysis. a, Influence of experiment domain on difference in pseudo-R2 values. b, Influence of prompt properties on difference in pseudo-R2 values. Positive values for a given variable indicate that finetuning led to an improved goodness-of-fit. Error bars depict the standard error.
Benchmarks
metabench
Figure 6: metabench [30] results.
CogBench
Newell test
Together with his call for unified theories of cognition [2], Newell outlined a set of criteria that a unified computational model should fulfill. In the following, we discuss Centaur in the light of a modified version proposed by Anderson and Lebiere [36].
@@ Acknowledgements
Psych-101 data
Refer to caption Refer to caption
The following sections present the included experiments in detail. Example prompts are truncated to $4096$ characters but otherwise shown as is.

### v3

@@ Discussion
The present paper introduced Centaur, the first foundation model of human cognition. We obtained Centaur by finetuning a state-of-the-art language model on Psych-101 – a novel, large-scale data set of human behavior. This approach allowed us to leverage the vast knowledge embedded in large language models, while, at the same time, aligning them with human behavior [15]. Centaur successfully captures human behavior and passes a wide range of out-of-distribution checks. It generalizes not only to unseen participants but also to new cover stories, structural variations, and entirely novel domains. In addition to analyzing the model on a behavioral level, we also conducted a series of analyses on its internal representations, in which we found an increased alignment with human neural activity.
We furthermore conducted a case study demonstrating how both Psych-101 and Centaur can be used for guiding the development of predictive, yet interpretable, cognitive models. Looking beyond this example, Centaur finds many additional applications in the context of automated cognitive science [49, 50, 51]. It may, for instance, be used for in-silico prototyping of experimental studies [52]. In this context, one could use the model to figure out which designs lead to the largest effect sizes, how to design a study to reduce the number of required participants, or to estimate the power of an effect.
While the present paper takes initial steps in leveraging Centaur to gain deeper insights into human cognition, it also opens up exciting new avenues for future exploration. First, one could further probe Centaur’s internal representations to understand how it represents knowledge and processes information. The resulting insights could, in turn, be used to generate new hypotheses about knowledge representation and information processing in humans that could be validated in future experimental studies. We believe that tools like sparse autoencoders [53] and attention map visualization [54] provide promising avenues towards accomplishing this goal and hope to explore them in future studies.
In addition, it might also be possible to train models with different architectures from scratch using the data set that we have created in the process of this paper. Doing so would enable us to investigate the neural architecture of human cognition at a scale that could not have been done before. We might, for example, ask questions such as whether human information processing is better described by attention-based architectures [55] or architectures with a vector-based memory [56], or how much we can improve by incorporating theories from the neuroscience literature [57, 58]. We expect an eventual outcome of such an approach to contain both domain-specific and domain-general modules, thereby allowing us to investigate the interplay between the two.
Even though Psych-101 is already the broadest and largest data set of human behavior out there, we view its development as an ongoing process and plan to further develop it. The focus in its current state is largely on learning and decision-making, but eventually, we intend to include additional domains such as psycholinguistics, social psychology, and economic games. Experiments with information about individual differences are another source of neglected data in the current iteration of Psych-101. Ideally, we want to include all types of relevant information about subjects (e.g., age, personality traits, or socioeconomic status) in the prompt, such that a model trained on this data can capture individual differences. Experiments from developmental psychology or computational psychiatry provide an ideal source for this purpose. Finally, while we have already included some cross-cultural and meta-studies [59, 60, 61, 62], the current iteration has still a strong bias toward a Western, educated, industrialized, rich, and democratic (WEIRD) population [63]. To address all of these shortcomings, we have created an open-source repository and invite everyone to contribute to the next iteration of Psych-101 in an open research collaboration (https://github.com/marcelbinz/Psych-201). The goal of this effort is to provide psychological data in a standardized format that facilitates benchmarking, thereby complementing existing efforts from the neuroscience community [64, 65, 66, 67]. While the natural language format (together with quite a bit of reverse-engineering) used in this work allows us to express a vast range of experimental paradigms, it introduces a selection bias against experiments that cannot be expressed in natural language. The long-term objective should be to move towards a multi-modal data format instead [68].
Conclusion
[…]
Methods
Data collection
Finetuning procedure
We used Llama 3.1 70B as the base model for our finetuning procedure. We relied on a parameter-efficient finetuning technique known as QLoRA [19]. QLoRA adds so-called low-rank adapters to each layer of a four-bit quantized base model. The base model was kept fixed during finetuning and only the parameters of the low-rank adapters were adjusted. We added low-rank adapters to all linear layers of the self-attention mechanisms and the feedforward networks. Each low-rank adapter modifies the forward pass as follows:
$\mathbf{Y}=\mathbf{X}\mathbf{W}+\alpha\mathbf{X}\mathbf{L}_{1}\mathbf{L}_{2}\qquad\qquad\mathbf{W}\in\mathbb{R}^{h\times o};~{}\mathbf{L}_{1}\in\mathbb{R}^{h\times r};~{}\mathbf{L}_{2}\in\mathbb{R}^{r\times o}$
where $\mathbf{X}\mathbf{W}$ is the (quantized) linear transformation of the base model and $\mathbf{X}\mathbf{L}_{1}\mathbf{L}_{2}$ the low-rank adapter. The hyperparameter $\alpha$ controls the trade-off between the two. Low-rank adapter computations are performed in half-precision floating-point format. For further details on this technique, we refer the reader to [19].
We finetuned the model for one epoch on the entire data set using a standard cross-entropy loss (we experimented with prolonged training but found that this led to overfitting). We only backpropagated the loss at human responses and masked out the loss for all other tokens. The effective batch size was set to 32, the learning rate to 0.00005, and the weight decay to 0.01. We used an 8-bit AdamW optimizer [70] with a linearly increasing warm up over the first 100 gradient steps. The finetuning procedure was implemented using the unsloth library (https://unsloth.ai/).
We have also trained a smaller version of Centaur – called Minitaur – that uses Llama 3.1 8B as base model following the same recipe. Minitaur captures human behavior close to its training distribution but generalizes less robustly than the larger model to out-of-distribution experiments. Nevertheless, we believe that Mintaur is useful for prototyping as it does not require access to any specific hardware (it runs, for instance, on the free GPU instances in Google Colab). More detailed results for this model are provided in the Supplementary Information.
Evaluation metric
We use (negative) log-likelihoods averaged over responses as our evaluation metric. For experiments with multi-token responses, we summed log-likelihoods within a response and averaged across responses.
Domain-specific cognitive models
For our main analysis, we were interested in predicting the behavior of held-out participants. Therefore, we fitted a joint set of parameters for all participants in the training data and evaluated how well a model with these parameters predicts responses of held-out participants. Mirroring the evaluation metric for the language-based models, we evaluated goodness-of-fit using (negative) log-likelihoods averaged over responses.
Neural alignment
The neural alignment analysis on the two-step task was conducted using data collected in a previous study [41]. We used a regularized linear regression model to predict fMRI data from internal representations of Centaur and Llama (a separate model was used for each participant and region). We fitted each of these models on data from two scanning blocks and evaluated them on data from the third. The regularization strength was selected using a nested cross-validation procedure. For each run, we split the beta maps into cortical and subcortical regions of interest (ROI) using the Schaefer 2018 atlas [71]. We averaged the betas within each ROI, reducing the number of betas from the number of voxels to the number of ROIs. All cortical and subcortical ROIs from the atlas were evaluated. Reported Pearson correlation coefficients correspond to the average across all ROIs.
Model-guided scientific discovery
In our model-guided scientific discovery analysis, we focused on participants in the test set to avoid any potential contamination issues. We fitted parameters of all cognitive models individually for each participant using a maximum likelihood estimation. Models were compared against each other using the Akaike information criterion (AIC). The three models from the original study were implemented via the following equations:
$p(a=\text{A}|\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}},\text{WADD})$ $\propto\exp\left(\beta\cdot\mathbf{w}_{\text{WADD}}^{\top}\mathbf{x}_{\text{A}}\right)$
$\mathbf{w}_{\text{WADD}}$ $=\left[0.9,0.8,0.7,0.6\right]$
$p(a=\text{A}|\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}},\text{EW})$ $\propto\exp\left(\beta\cdot\mathbf{w}_{\text{EW}}^{\top}\mathbf{x}_{\text{A}}\right)$
$\mathbf{w}_{\text{EW}}$ $=\left[1,1,1,1\right]$
$p(a=\text{A}|\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}},\text{TTB})$ $\propto\exp\left(\beta\cdot\mathbf{w}_{\text{TTB}}^{\top}\mathbf{x}_{\text{A}}\right)$
$\mathbf{w}_{\text{TTB}}$ $=\left[1,0.5,0.25,0.125\right]$
where $\mathbf{x}_{\text{A}}$ and $\mathbf{x}_{\text{B}}$ are vectors containing four expert ratings (either 0 or 1) and $\beta$ is a free parameter controlling the noise level.
We prompted DeepSeek-R1 (in the Distill-Llama-70B variant) to generate explanations of human decision-making; the corresponding prompt is provided in the Supplementary Information. We then formalized the explanation shown in Figure 5 into the following computational model:
[…]
For the scientific regret minimization pipeline, we computed the difference in log-likelihoods between Centaur and the DeepSeek-R1-discovered model. We visualized and inspected the ten data points with the highest difference. This process resulted in the following computational model:
$p(a=\text{A}|\mathbf{x}_{\text{A}},\mathbf{x}_{\text{B}},\text{SRM})$ $\propto\exp\left(\beta\cdot\left(\sigma\cdot\mathbf{w}_{\text{TTB}}^{\top}\mathbf{x}_{\text{A}}+\left(1-\sigma\right)\cdot\mathbf{w}_{\text{EW}}^{\top}\mathbf{x}_{\text{A}}\right)\right)$
where $\sigma$ is a free parameter constrained between 00 and $1$ controlling the trade-off between the two strategies.
Data availability
Code availability
[…]
Acknowledgements
[…]
Author contributions
Project lead: Marcel Binz
Data curation: Elif Akata, Franziska Brändle, Marcel Binz, Fred Callaway, Julian Coda-Forno, Can Demircan, Maria Eckstein, Noemi Elteto, Susanne Haridi, Akshay Jagadish, Li Ji-An, Alexander Kipnis, Sreejan Kumar, Tobias Ludwig, Surabhi Nath, Joshua Peterson, Evan Russek, Tankred Saanum, Johannes Schubert, Luca Schulze Buschoff, Nishad Singhi, Xin Sui, Mirko Thalmann, Vuong Truong, Kristin Witte, Shuchen Wu, Dirk Wulff, Huadong Xiong
Model training: Marcel Binz, Vishaal Udandarao
Model evaluation: Marcel Binz, Julian Coda-Forno, Alexander Kipnis, Mirko Thalmann, Konstantinos Voudouris
First draft: Marcel Binz, Eric Schulz
Review and editing: All authors
No researchers at Google DeepMind used Llama for this research. We thank Natalia Scharfenberg for her contributions to the data collection.
Supplementary Information
Psych-101
Refer to caption Refer to caption
Data contamination analysis
We performed a data contamination analysis using the LogProber method proposed in [76]. LogProber fits the following two-parameter exponential model to the cumulative log-likelihood of each sequence being checked for contamination:
$f(x)=-A\left(1-e^{-Bx}\right)$
Prompts that are memorized from the pretraining data will show a high acceleration ($B$). Following the results presented in [76], we set a threshold for possible contamination to $\log B\geq 1$. We repeat this analysis for each experiment in Psych-101. For this, we used prompts up to the point of the first human choice. This analysis indicated no evidence of contamination as shown in Figure 7.
Figure 7: Data contamination analysis using the LogProber method [76]. LogProber fits an exponential model to the cumulative log-likelihood of each sequence being checked for contamination. High acceleration ($\log B$) suggests that a prompt is memorized from the pretraining data.
Experiment embeddings
Figure 8 shows a two-dimensional embedding of the experiments used in this paper. To obtain this embedding, we took the corresponding natural language prompts up to the point of the first human choice, extracted a vector-based representation for them using ModernBERT [77], and finally projected these representations onto two dimensions using multidimensional scaling. We observed that most of our out-of-distribution evaluation experiments are placed on the boundaries of the embedding space, confirming that they indeed probe the generalization abilities of Centaur.
Figure 8: Multidimensional scaling embedding of the experiments used in this paper. Purple dots correspond to experiments from Psych-101, whereas the colored dots correspond to the indicated evaluation experiment.
Full goodness-of-fit results
Experiment Centaur Llama Cognitive model
Shepard categorization 0.5394 0.5818 0.6108
Drifting four-armed bandit 0.7029 0.8810 0.9043
N-back 0.3954 0.5209 0.5787
Digit span 0.5520 0.6618 0.9359
Go/no-go 0.0000 0.0062 0.0757
Recent probes 0.2572 0.3433 0.3868
Horizon task 0.4032 0.5237 0.3595
Gardening task 0.3783 0.5040 0.9105
Columbia card task 0.1867 0.2261 0.2629
Balloon analog risk task 0.0593 0.0753 0.0922
Two-armed bandit 0.2963 0.3829 0.4187
Conditional associative learning 0.5380 0.6373 0.8575
THINGS odd-one-out 0.8068 1.1386 0.8253
Multi-attribute decision-making 0.0619 0.1502 0.1922
Two-step task 0.4998 0.6075 0.6043
Probabilistic instrumental learning 0.4937 0.5382 0.5047
Medin categorization 0.4967 0.5772 0.5313
Zoopermarket 0.4850 0.6026 0.6047
choices13k 0.4274 0.5342 0.6563
CPC18 0.3390 0.4118 0.6607
Intertemporal choice 0.4340 0.7336 0.6591
Structured bandit 0.6410 0.8114 1.0530
Weather prediction task 0.5514 0.5749 0.6267
Iowa gambling task 0.8890 0.9880 1.1555
Virtual subway network 1.1271 1.5347 nan
Multi-task reinforcement learning 0.5672 0.6604 1.0424
Serial reaction time task 0.1718 0.1900 0.1962
Decisions from description 0.5336 0.7569 0.6120
Decisions from experience 0.3686 0.4339 0.5404
Changing bandit 0.3025 0.3824 0.4378
Multiple-cue judgment 1.1236 1.2818 1.9157
Recall and recognition 1.0591 1.3759 nan
Experiential-symbolic task 0.4536 0.6983 nan
Grammar judgement 1.4355 1.9949 1.4127
Risky choice 0.4281 0.6475 nan
Tile-revealing task 1.8713 2.7380 nan
Episodic long-term memory 0.8684 1.1344 nan
Aversive learning 4.0733 5.1066 nan
Spatially correlated multi-armed bandit 1.8319 2.4479 2.7635
Probabilistic reasoning 2.3731 2.6406 nan
Table 1: Full negative log-likelihoods on held-out participants.
Minitaur
Figure 9 compares Centaur and to Minitaur on the analyses from the main text.
Figure 9: Log-likelihood comparison between Centaur and Minitaur on the analyses from the main text. The left subplot contains results from Psych-101, while the right subplot shows contains results for different held-out settings. Error bars correspond to the standard error of the mean, taken over responses.
Further out-of-distribution evaluations
Figure 10 contains results for six additional out-of-distribution experimental paradigms, including moral decision-making [31], economic games [32], naturalistic category and reward learning [33], behavioral propensities [34], and a deep sequential decision task [35]. None of these paradigms were included in Psych-101, hence they provide a stress test for a model’s generalization capabilities. Centaur robustly captured human behavior in all of these settings, while smaller and non-finetuned models did not do so consistently.
Figure 10: Further out-of-distribution evaluations. Each subplot shows negative log-likelihoods for a different experiment. Error bars correspond to the standard error of the mean, taken over responses.
Further model comparisons
To rule out the hypothesis that finetuning on any data aligns a model with human behavior, we compared Centaur to various Llama variants finetuned for other purposes (i.e. non-cognitive tasks). These models include the following:
  • Nemotron [78]: finetuned for instruction-following.
  • Hermes [79]: finetuned for various purposes, including agentic capabilities, roleplaying, reasoning, multi-turn conversation, and long context coherence.
  • Reflection: finetuned for reasoning.
Figure 11 illustrates the results of this analysis. We observed that none of the Llama variants captures human behavior better than the base model, ruling out the hypothesis that finetuning generally leads to models that are better at predicting human behavior.
Figure 11: Negative log-likelihoods of Centaur and alternative Llama variants on Psych-101. Error bars correspond to the standard error of the mean, taken over responses.
Noise ceiling analysis
We conducted a noise ceiling analysis to better understand the capabilities of Centaur. It is not straightforward to estimate the noise ceiling for experiments with sequential dependencies, which includes the majority of Psych-101. Hence, we focused on two experiments for which such an analysis is possible: the choices13k data set [21] and an intertemporal choice experiment [59]. In both cases, we found that Centaur substantially exceeds the estimated noise ceiling. This is possible because Centaur can pick up on context-dependent patterns that are not captured by standard noise ceiling analyses. Therefore, we have performed an additional analysis testing how well Centaur can predict human responses if we prompt it to predict each response independently. We use the suffix “ind.” to indicate this way of prompting the model. Centaur still matches the performance of domain-specific cognitive models when context-independent prompts are used, amounting to roughly half of the estimated noise ceiling. Figure 12 visualizes these results.
Figure 12: Noise ceiling analysis. We use the suffix ”ind.” to indicate models that are prompted with to predict each response independently.
Benchmarks
metabench
Figure 13: metabench [38] results.
CogBench
Fine-grained neural analysis results
Figure 15 shows fine-grained neural alignment results on the two-step task data. Centaur achieves the most accurate predictions in the left motor cortex (see Figure 15a). As participants performed the task with their right hand in the scanner, this effect may be explained by Centaur’s strong performance in predicting choices.
Figure 15b contains numerical results for ROIs that have been identified as behaviorally relevant in previous work. Centaur outperformed Llama and the cognitive model in predicting activity in accumbens, the ROI from the original study that showed a reward prediction error effect [22, 41]. We found a similar pattern in the medial PFC, another region that showed an effect in the original article [41], as well as in the sensory and motor cortices.
Figure 15: Neural alignment results in the two-step task. a, Pearson correlation coefficients between the predicted activity from Centaur’s representations and the BOLD data shown on a surface brain. b, Predictive performance of Centaur’s representations against alternatives for specific ROIs. Cortical scores are averaged over the corresponding bilateral parcels in the Schaefer atlas. The accumbens is defined based on the Harvard-Oxford atlas. Pearson correlation coefficients are shown for layer 20 but exhibit a similar pattern across all layers.
Model-guided scientific discovery
We generated an explanation for the behavior of every participant in the multi-attribute decision-making study of [46] using DeepSeek-R1 based on the following prompt template:
[…]
Newell test
@@ Acknowledgements
Together with his call for unified theories of cognition [2, 81], Newell outlined a set of criteria that a unified computational model should fulfill. Centaur is the first model to satisfy the majority of these criteria (see Table 2). Most importantly, it (1) behaves as an almost arbitrary function of the environment, (2) operates in real time, and (3) relies on vast amounts of knowledge about the world. We provide an extended discussion on Newell’s criteria in the following.
@@
Table 2: Newell test for a theory of cognition.
@@
Example prompts
The following sections present the experiments contained in Psych-101 in detail. Example prompts are truncated to $4096$ characters but otherwise shown as is.